Given the table below, what is the probability of a number greater than 2 and less than 5?

Simple Event

Probability

1

0.20

2

0.12

3

0.16

4

0.18

5

0.15

6

0.19
Answer #1

From given probability distribution ,

P(x > 2 and x < 5)

= P(x = 3) + P(x = 4)

= 0.16 + 0.18

= 0.34

Probability = 0.34